x+10-6x=-5(x+2)
-5x +10=-5x+10
0=0
Is this correct

Answer by Earlsdon(6294)

No, it is not!
If you start with:
x+10-6x = -5(x+2) and simplify, you'll get:
10-5x = -5x-10 or
-5x + 10 = -5x - 10 which cannot be true no matter what value of x you use.
-5x + 10 = -5x - 10 Add 5x to both sides.
10 = -10 No way!